OpenStudy (anonymous):

state if this has no solution,and set in builder notation 3x + 4y = 5 2x + y = 1

OpenStudy (amistre64):

in order to be no solution they have to be the same equation and differ only by a constant; and maybe by scale

OpenStudy (anonymous):

like same slope

OpenStudy (anonymous):

x=(5*1-4)/(3*1-2*4)=-1/5 y=(3-10)/(3*1-2*4)=7/5 may b...

OpenStudy (anonymous):

their slopes are different so ur good. y=-3/4 x+5 y=-2x+1

OpenStudy (amistre64):

if they want you to find a solution; either sub it or eliminate it; id go with elimination since it just sounds dirty lol